# MMEDIT: A Unified Framework for Multi-Type Audio Editing via Audio Language Model

### v1

@@ Abstract
Text-guided audio editing requires modifying specific acoustic events while strictly preserving non-target content. Despite recent progress, existing approaches remain fundamentally limited: training-free methods often suffer from signal degradation due to diffusion inversion. Training-based methods exhibit better generation quality but remain severely constrained by the scarcity of high-quality paired data and task formulations that cover only a narrow subset of editing operations. Moreover, standard architectures in this paradigm often decouple text and audio processing, limiting the model’s ability to align instructions with specific acoustic contexts. To overcome these constraints, we propose MMEdit, an ALM-driven framework for unified audio editing. We systematically extend task definitions to cover a comprehensive range of operations, including addition, replacement, removal, reordering, and attribute modification. Furthermore, we construct a scalable synthesis pipeline to generate large-scale paired data with fine-grained event-level annotations. To capture complex editing semantics, we integrate Qwen2-Audio encoder with an MMDiT-based generator, achieving precise cross-modal alignment and localized editing. Experiments demonstrate that our approach delivers superior editing localization accuracy, robust instruction following, and high fidelity in non-edited regions. The project page can be accessed at Project Page.
@@ I Introduction
With the rapid development of generative modeling, modern text-to-audio (TTA) systems [yang2023diffsound, liu2023audioldm, majumder2024tango] are now capable of producing high-fidelity, diverse, and text-aligned sounds from natural language prompts. Unlike TTA, which synthesizes audio from scratch, text-guided audio editing seeks to modify specific regions of an existing audio signal according to user-provided textual instructions, while preserving the remaining content unchanged. This paradigm offers users a more intuitive and efficient way to achieve fine-grained control over audio content through natural language, rather than relying on manual waveform editing or expert post-processing. Since the introduction of this task in AUDIT [wang2023audit], research in this field has attracted increasing research interest [xu2024prompt, manor2024zero, jia2025audioeditor, paissan2024audio].
@@
Existing studies on text-guided audio editing can be broadly categorized into training-free and training-based paradigms. Training-free approaches typically adapt pre-trained diffusion-based TTA models by inverting the denoising process to reconstruct latent noise and re-generate audio under new textual conditions. PPAE [xu2024prompt] and AudioEditor [jia2025audioeditor] are representative methods. Despite free from additional training, these approaches rely heavily on diffusion inversion, resulting in high inference cost and a strong dependence on full captions or handcrafted masks, restricting their applicability to complex editing tasks. Crucially, the mechanism of inversion and subsequent regeneration may compromise the fidelity to the original structure and fine details.
@@
In contrast, training-based frameworks explicitly learn instruction-conditioned generation models from paired data. AUDIT [wang2023audit] pioneered this paradigm by training a latent diffusion model. Paissan et al. [paissan2024audio] explored prompt interpolation and fine-tuning within the embedding space, while RFM-Editing [gao2025rfm] incorporated rectified flow matching to improve efficiency. Compared to training-free inversion, this paradigm offers better fidelity and faster inference speeds.
@@
Despite these advances, the progress of audio editing faces challenges in both data resources and model design. In terms of data, real editing pairs are rarely available, and existing studies thus rely on small-scale or manually synthesized datasets to simulate editing operations. However, prior synthetic pipelines often adopt simple mixing strategies with limited event-level controllability, resulting in incomplete coverage of edit types, and therefore failing to reflect a comprehensive range of real-world manipulations. The absence of an open and unified benchmark further exacerbates the challenge, leading to inconsistent evaluation and limited reproducibility across different studies. In terms of model design, most existing works rely on text-only encoders to process instructions. The interaction between textual instructions and the source audio is limited to the cross-attention mechanism in the generation model. Yet in complex audio scenes, accurately editing a specific event entails not only interpreting the instruction but also locating the event within the source audio, a capability that text-only semantic encoders inherently lack. Consequently, such a decoupled, late-fusion design is often inadequate for achieving context-aware understanding of complex editing instructions.
@@
To address these limitations, we propose MMEdit, a unified framework driven by an Audio Language Model for general, instruction-following audio editing. First, we systematically expand the editing scope to six representative operations: addition, removal, replacement, loudness adjustment, reordering, and speed modification, and develop a scalable synthetic pipeline, together with a unified benchmark for standardized evaluation. Although some operations can be achieved by signal processing tools in single-event audio, event overlap in realistic soundscapes makes event-specific editing difficult. Therefore we integrate them into a unified editing model to learn shared audio editing capabilities across operations. Moreover, MMEdit utilizes an Audio Language Model (ALM) to process instructions and audio content jointly. This unified processing enables precise cross-modal alignment, allowing the model to precisely interpret instructions within the specific acoustic context. The ALM encoder is integrated with an MMDiT-based generator [esser2024scaling] to generate high-fidelity audio. Our main contributions are summarized as follows:
@@
  • Scalable Data Pipeline and Unified Benchmark: We develop a synthetic generation pipeline that produces large-scale paired data with fine-grained, event-level annotations. To foster reproducibility, we establish and open-source an evaluation benchmark.  [^1] Will be released upon acceptance.
@@ II-A Audio Segment Curation
Existing text-guided audio editing studies like AUDIT typically synthesize training pairs by combining entire audio clips with their full global captions. However, to enable more precise and interpretable editing such as specifying event timing and occurrence counts, a finer granularity is required. We first perform event-level segmentation and grounding on an audio-caption corpora, AudioCaps [kim2019audiocaps]. Specifically, inspired by AudioTime [xie2025audiotime], we employ an LLM to decompose each caption into single-event descriptions. The descriptions are then timestamped by the Text-to-Audio Grounding (TAG) model [xu2024towards] to determine event boundaries, and only the non-overlapping segments are retained and cropped as base audio units for subsequent synthesis. Finally, to guarantee semantic alignment, we compute CLAP scores [elizalde2023clap] between the textual label and the extracted segment and discard pairs with a similarity score below $0.3$.
@@ II-B Editing Task Definition and Composition
Building upon the disentangled foreground and background components, we systematically define six representative editing operations that simulate real-world acoustic manipulations. We formulate each operation as a transformation function applied to the foreground and background streams, parameterized by key attributes: event type, quantity, and temporal position. To technically implement these compositional processes, we leverage the Scaper toolkit [salamon2017scaper] for programmable soundscape synthesis and augmentation. This allows for precise control over signal-to-noise ratios (SNR), event timing, and dynamic mixing. This unified formulation accommodates multiple subtypes within each category, enabling a flexible and scalable framework for diverse audio editing scenarios.
@@ II-B6 Speed Modification
Given the structured nature of audio editing instructions, we designed diverse prompt templates for each editing type, incorporating specific slots for variable parameters such as event types, timestamps, and occurrence counts. We then construct corresponding textual instructions by populating these slots with fine-grained metadata to simulate realistic user commands. In total, the simulated corpus comprises over one million triplets, amounting to approximately 2500 hours of audio. Furthermore, to ensure a reliable evaluation benchmark, we perform an additional rigorous screening process to filter out a test set, with high audio fidelity and semantic alignment guaranteed. We make both this benchmark and our generation pipeline publicly available to foster reproducible research and facilitate fair comparison. The detailed statistics for each editing task are presented in Table I.
TABLE I: Statistics of MMEdit Dataset. Number of samples for each editing task.
@@ III MMEDIT
In this section, we introduce MMEdit, a general-purpose audio editing framework capable of handling diverse editing tasks, which integrates an audio LLM for understanding and MMDiT [esser2024scaling] for high-quality audio generation. As illustrated in Fig. 2, MMEdit consists of three major components: a waveform variational autoencoder (VAE) [xu2025uniflow] for latent audio representation, an Audio Language Model (ALM)-based encoder for cross-modal understanding, and an MMDiT-based backbone that performs generation in the latent space.
@@ III-A Multimodal Understanding
Crucial to general editing capabilities is the model’s ability to deeply understand the interaction between the source audio content and the textual instruction. Instead of using a text-only encoder (e.g., T5) for instruction understanding, we use Qwen2-Audio [chu2024qwen2] as a multimodal encoder to obtain a more representative semantic feature to guide editing. Compared with text-only encoders, Qwen2-Audio can jointly process audio and text inputs, yielding inherently cross-modal aligned representations. Formally, given a source audio clip $\mathbf{x}_{in}$ and a textual instruction $\mathbf{y}$, the ALM produces a sequence of embeddings:
@@ III-B1 Global Conditioning via Adaptive Modulation
The global context $\mathbf{c}_{global}$ from the ALM is fused with the diffusion timestep $\mathbf{t}$. The resulting vector modulates the entire network through adaptive Layer Normalization (AdaLN) [peebles2023scalable] before each Transformer block, serving as a global semantic guidance for generation.
@@ III-B2 MMDiT Blocks Design
The backbone comprises a stack of $N_{\text{joint}}$ Joint Blocks followed by $N_{\text{single}}$ Single Blocks. In each joint block, the audio latents and the multimodal instruction embeddings are concatenated along the sequence dimension and processed by a shared self-attention layer. This joint attention mechanism lets latent audio directly attend to instruction embeddings, enabling event-level, text-guided editing. After joint blocks, single blocks operate purely within audio latents using self-attention. Single blocks focus on refining temporal continuity and spectral coherence, ensuring that the edited audio remains acoustically natural and that unedited regions are preserved with high fidelity.
@@ III-C Latent Diffusion for Audio Editing
During training, we follow the standard DDPM [ho2020denoising] formulation. For a randomly sampled timestep $t\in\{1,\dots,T\}$, a noisy version $\mathbf{z}_{t}$ of the target latent is obtained by
@@
During training, text instructions are randomly masked to learn unconditional generation while source audio latent is always fed to the model. During inference, we use DDIM [songdenoising] sampler with classifier-free guidance, where the guided noise estimate is formulated as $\tilde{\epsilon}=\epsilon_{\text{uncond}}+w(\epsilon_{\text{cond}}-\epsilon_{\text{uncond}}).$ Finally, the VAE decoder converts the denoised latent into the edited audio.
@@ IV-A Training Setup
We compare MMEdit with two representative audio editing systems: AUDIT and AudioEditor. Both only support Addition, Removal, and Replacement operations. Since AudioEditor operates on source–target caption pairs, we rewrite each editing instruction into a full-sentence target caption to align with its required input format. To comprehensively assess performance, we report both reference-based objective metrics on synthetic pairs and Mean Opinion Score (MOS)-based subjective metrics on real recordings.
@@ IV-B1 Objective evaluation
On the synthetic test set, we report log-spectral distance (LSD), Fréchet Audio Distance (FAD), Fréchet Distance (FD), Kullback–Leibler (KL) divergence, and Inception Score (IS) [liu2023audioldm] as our main metrics. LSD, FAD, FD, and KL measure spectral or distributional discrepancies between the generated audio and the reference target in pretrained embedding spaces, where smaller values indicate closer alignment to the ground-truth. IS reflects the perceptual quality and diversity of the generated audio, with larger values corresponding to more realistic and varied samples.
@@ IV-B2 Subjective evaluation
To mitigate the potential bias inherent in synthetic data, we conduct a listening test on a curated subset of real AudioCaps recordings paired with manually designed instructions. As no ground-truth edited audio is available, we rely on human judgment. For all tasks, we invite ten college-level raters to do a subjective rating with explicit instructions. Each edited sample is rated on a 1–5 Likert scale along two dimensions: Relevance (R-MOS), which measures how accurately the system follows the editing instruction; and Faithfulness (F-MOS), which measures how well non-edited regions of the original audio are preserved without introducing unintended artifacts or distortions.
@@
Reorder MMEdit 2.353 0.791 9.328 0.391 5.276 3.90 4.40
Loudness MMEdit 1.310 0.844 15.869 0.597 4.864 4.43 4.35
@@ IV-C Main Results and Analysis
Consistent trends are observed on real AudioCaps clips under the subjective listening test. As detailed in Table II, MMEdit achieves the highest scores across both metrics. In terms of R-MOS, MMEdit consistently scores above 3.7, indicating strong instruction following. The advantage is even more pronounced in F-MOS, where MMEdit demonstrates superior preservation of non-edited regions. Notably, despite high IS, AudioEditor receives markedly lower subjective scores, suggesting more aggressive modifications to maximize diversity at the cost of destroying non-editing contents.
@@ IV-D Ablation Studies
We investigate individual contributions of the ALM encoder, the diffusion backbone architecture, and the classifier-free guidance strategy [ho2021classifier].
@@ IV-D1 Effect of ALM Encoder
To assess the critical role of the multimodal ALM, we conduct an ablation study by replacing the Qwen2-Audio encoder with a text-only encoder (Flan-T5 [chung2024scaling]), keeping other components invariant. As presented in Table III, although the text-only variant (w/ T5) achieves higher IS, the decline in fidelity metrics indicates that the text-only encoder fails to capture the intricate acoustic cues of the source audio. In contrast, Qwen2-Audio contributes richer cross-modal information by jointly encoding the instruction and the source audio, enabling more accurate and perceptually coherent edits than a text-only encoder could provide.
TABLE III: Ablation study on model components. Note that subjective evaluation was omitted for w/ DiT as it underperforms the full model on all objective metrics.

### v2

@@ Abstract
Text-guided audio editing requires modifying specific acoustic events while strictly preserving non-target content. Despite recent progress, existing approaches remain fundamentally limited: most are restricted to elementary operations (e.g., addition and removal), while real-world scenarios require more extensive tasks. Moreover, standard architectures in this paradigm often decouple text and audio processing, limiting the model’s ability to align instructions with specific acoustic contexts. To overcome these constraints, we propose MMEdit, an ALM-driven framework for unified audio editing. We systematically extend task definitions to cover a comprehensive range of operations, including addition, replacement, removal, reordering, and attribute modification. Furthermore, we construct a scalable synthesis pipeline to generate large-scale paired data with fine-grained event-level annotations. To capture complex editing semantics, we integrate Qwen2-Audio encoder with an MMDiT-based generator, achieving precise cross-modal alignment and localized editing. Experiments demonstrate that our approach delivers superior editing localization accuracy, robust instruction following, and high fidelity in non-edited regions. The project page can be accessed at Project Page.
@@ I Introduction
With the rapid development of generative modeling, modern text-to-audio (TTA) systems [22, 10, 11] are now capable of producing high-fidelity, diverse, and text-aligned sounds from natural language prompts. Unlike TTA, which synthesizes audio from scratch, text-guided audio editing (TTE) seeks to modify specific regions of an existing audio signal according to user-provided textual instructions, while preserving the remaining content unchanged [17]. This paradigm offers users a more intuitive and efficient way to achieve fine-grained control over audio content through natural language, rather than relying on manual waveform editing or expert post-processing. Since the introduction of this task in AUDIT [17], audio editing has attracted increasing research interest [19, 12, 8, 13].
@@
Existing studies on TTE can be broadly categorized into training-free and training-based paradigms. Training-free approaches typically adapt pre-trained diffusion-based TTA models by inverting the denoising process to reconstruct latent noise and re-generate audio under new textual conditions. PPAE [19] and AudioEditor [8] are representative methods. Although free from additional training, these approaches rely heavily on diffusion inversion, resulting in high inference cost and a strong dependence on full captions or handcrafted masks, restricting their applicability to complex editing tasks. Crucially, the mechanism of inversion and subsequent regeneration may compromise the fidelity to the original structure and fine details.
@@
In contrast, training-based frameworks explicitly learn instruction-conditioned generation models from paired data. AUDIT [17] pioneered this paradigm by training a latent diffusion model. Paissan et al. [13] explored prompt interpolation and fine-tuning within the embedding space, while RFM-Editing [5] incorporated rectified flow matching to improve efficiency. Compared to training-free inversion, this paradigm offers better fidelity and faster inference speeds.
@@
Despite these advances, the progress of audio editing faces challenges in both data resources and model design. In terms of data, real editing pairs are rarely available, hence existing studies rely on small-scale or manually synthesized datasets to simulate editing operations[17, 8]. However, prior synthetic pipelines often adopt simple mixing strategies with limited event-level controllability, resulting in coverage of only basic edit types (i.e., addition and removal), and therefore failing to reflect a comprehensive range of real-world manipulations. The absence of an open and unified benchmark further exacerbates the challenge, leading to inconsistent evaluation and limited reproducibility across different studies. In terms of model design, most existing works rely on text-only encoders to process instructions. The interaction between textual instructions and the source audio is limited to the cross-attention mechanism in the generation model. Yet in complex audio scenes, accurately editing a specific event entails not only interpreting the instruction but also locating the event within the source audio, a capability that text-only semantic encoders inherently lack. Consequently, such a decoupled, late-fusion design is often inadequate for achieving context-aware understanding of complex editing instructions.
@@
To address these limitations, we propose MMEdit, a unified framework driven by an Audio Language Model for general, instruction-following audio editing. First, we systematically expand the editing scope to six representative operations: addition, removal, replacement, loudness adjustment, reordering, and speed modification for better adaptation with real-world applications. We thereby develop a scalable pipeline yielding over one million training samples, distilling a high-quality, human-verified benchmark for evaluation. Although some operations can be achieved by signal processing tools in single-event audio, event overlap in realistic soundscapes makes event-specific editing difficult. Therefore we integrate them into a unified editing model to learn shared audio editing capabilities across operations. Moreover, MMEdit utilizes an Audio Language Model (ALM) to process instructions and audio content jointly. This unified processing enables precise cross-modal alignment, allowing the model to precisely interpret instructions within the specific acoustic context. The ALM encoder is integrated with an MMDiT-based generator [4] to generate high-fidelity audio. Our main contributions are summarized as follows:
@@
  • Scalable Data Pipeline and Unified Benchmark: We develop a synthetic generation pipeline that produces large-scale paired data with fine-grained, event-level annotations. To foster reproducibility, we establish and open-source an evaluation benchmark on TTE for the first time.
@@ II-A Audio Segment Curation
Existing TTE studies like AUDIT typically synthesize training pairs by combining entire audio clips with their full global captions. However, to enable more precise and interpretable editing such as specifying event timing and occurrence counts, a finer granularity is required. We first perform event-level segmentation and grounding on an audio-caption corpus, AudioCaps [9]. Specifically, inspired by AudioTime [18], we employ an LLM to decompose each caption into single-event descriptions. The descriptions are then timestamped by the Text-to-Audio Grounding (TAG) model [20] to determine event boundaries, and only the non-overlapping segments are retained and cropped as base audio units for subsequent synthesis. Finally, to guarantee semantic alignment, we compute CLAP scores [3] between the textual label and the extracted segment and discard pairs with a similarity score below $0.3$.
@@ II-B Editing Task Definition and Composition
Building upon the disentangled foreground and background components, we systematically define six representative editing operations that simulate real-world acoustic manipulations. We formulate each operation as a transformation function applied to the foreground and background streams, parameterized by key attributes: event type, quantity, and temporal position. In our notation, $F_{n}$ denotes the $n$-th foreground event, while $F^{\prime}$ and $B^{\prime}$ represent new foreground and background components, respectively. To technically implement these compositional processes, we leverage the Scaper toolkit [15] for programmable soundscape synthesis and augmentation. This allows for precise control over signal-to-noise ratios (SNR), event timing, and dynamic mixing. This unified formulation accommodates multiple subtypes within each category, enabling a flexible and scalable framework for diverse audio editing scenarios.
@@ II-B6 Speed Modification
Given the structured nature of audio editing instructions, we designed diverse prompt templates for each editing type, incorporating specific slots for variable parameters such as event types, timestamps, and occurrence counts. We then construct corresponding textual instructions by populating these slots with fine-grained metadata to simulate realistic user commands. By scaling this pipeline, we construct a massive training corpus comprising over one million triplets, amounting to approximately 2500 hours of audio. To ensure a reliable evaluation standard, we further curate a high-quality benchmark through rigorous screening and human verification. We make both this benchmark and our generation pipeline publicly available to foster reproducible research and facilitate fair comparison. The detailed statistics for each editing task are presented in Table I.
TABLE I: Statistics of MMEdit Data. The distribution of training samples and the benchmark across editing tasks.
@@ III MMEDIT
In this section, we introduce MMEdit, a general-purpose framework designed for instruction-following audio editing. Distinct from prior decoupled designs that rely on cross-attention to bridge modalities, MMEdit integrates an ALM to process instructions and audio jointly for deep alignment, combined with an MMDiT backbone [4] for high-fidelity generation. This design enables more precise cross-modal alignment, allowing the model to interpret instructions within the specific acoustic context. As illustrated in Fig. 2, MMEdit consists of three major components: a waveform variational autoencoder (VAE) [21] for latent audio representation, an Audio Language Model (ALM)-based encoder for cross-modal understanding, and an MMDiT-based backbone that performs generation in the latent space.
@@ III-A Multimodal Understanding
Crucial to general editing capabilities is the model’s ability to deeply understand the interaction between the source audio content and the textual instruction. Instead of relying on a text-only encoder (e.g., T5) for instruction understanding, we use Qwen2-Audio [1] as a multimodal encoder to obtain a more representative semantic feature to guide editing. Qwen2-Audio can jointly process the source audio $\mathbf{x}_{in}$ and instruction $\mathbf{y}$, yielding inherently cross-modal aligned representations:
@@ III-B1 Global Conditioning via Adaptive Modulation
The global context $\mathbf{c}_{global}$ from the ALM is fused with the diffusion timestep $\mathbf{t}$. The resulting vector modulates the entire network through adaptive Layer Normalization (AdaLN) [14] before each Transformer block, serving as a global semantic guidance for generation.
@@ III-B2 MMDiT Blocks Design
The backbone comprises a stack of $N_{\text{joint}}$ Joint Blocks followed by $N_{\text{single}}$ Single Blocks. In each joint block, the audio latents and the multimodal instruction embeddings are concatenated along the sequence dimension and processed by a shared self-attention layer. This joint attention mechanism enables latent audio directly attend to instruction embeddings, enabling event-level, text-guided editing. After joint blocks, single blocks operate purely within audio latents using self-attention. Single blocks focus on refining temporal continuity and spectral coherence, ensuring that the edited audio remains acoustically natural and that unedited regions are preserved with high fidelity.
@@ III-C Latent Diffusion for Audio Editing
During training, we follow the standard DDPM [6] formulation. For a randomly sampled timestep $t\in\{1,\dots,T\}$, a noisy version $\mathbf{z}_{t}$ of the target latent is obtained by
@@
During training, text instructions are randomly masked to learn unconditional generation while source audio latent is always fed to the model. During inference, we use DDIM [16] sampler with classifier-free guidance, where the guided noise estimate is formulated as $\tilde{\epsilon}=\epsilon_{\text{uncond}}+w(\epsilon_{\text{cond}}-\epsilon_{\text{uncond}}).$ Finally, the VAE decoder converts the denoised latent into the edited audio.
@@ IV-A Training Setup
We compare MMEdit with two representative audio editing systems: AUDIT and AudioEditor, since they are the only publicly available models similar to our setting. Note that both only support Addition, Removal, and Replacement operations. Since AudioEditor operates on source–target caption pairs, we rewrite each editing instruction into a full-sentence target caption to align with its required input format. To comprehensively assess performance, we report both reference-based objective metrics on synthetic pairs and Mean Opinion Score (MOS)-based subjective metrics on real recordings.
@@ IV-B1 Objective evaluation
On the synthetic test set, we report Log-Spectral Distance (LSD), Fréchet Audio Distance (FAD), Fréchet Distance (FD), and KL divergence to measure spectral and distributional discrepancies, alongside Inception Score (IS) [10] for perceptual quality and diversity.
@@ IV-B2 Subjective evaluation
To mitigate the potential bias inherent in synthetic data, we conduct a listening test on a curated subset of 60 real AudioCaps recordings paired with manually designed instructions. As no ground-truth edited audio is available, we rely on human judgment. For all tasks, we invite ten college-level raters to do a subjective rating with explicit instructions. Each edited sample is rated on a 1–5 Likert scale along two dimensions: Relevance (R-MOS), which measures how accurately the system follows the editing instruction; and Faithfulness (F-MOS), which measures how well non-edited regions of the original audio are preserved without introducing unintended artifacts or distortions.
@@
Reorder MMEdit 2.353 0.791 9.328 0.391 6.381 3.90 4.40
Loudness MMEdit 1.310 0.447 8.528 0.252 6.321 4.43 4.35
@@ IV-C Main Results and Analysis
As detailed in Table II, subjective results on real AudioCaps clips also confirm MMEdit’s superiority. MMEdit consistently scores above 3.7 in R-MOS, indicating strong instruction following. This advantage is even more pronounced in F-MOS, where MMEdit demonstrates superior preservation of non-edited regions. The result further validated by paired $t$-tests ($p<0.001$) against both baselines. Notably, despite high IS, AudioEditor receives markedly lower subjective scores, suggesting more aggressive modifications to maximize diversity at the cost of destroying non-editing contents.
@@ IV-D Ablation Studies
We investigate individual contributions of the ALM encoder, the diffusion backbone architecture, and the classifier-free guidance strategy [7].
@@ IV-D1 Effect of ALM Encoder
To assess the critical role of the multimodal ALM, we conduct an ablation study by replacing the Qwen2-Audio encoder with a text-only encoder (Flan-T5 [2]), keeping other components invariant. As presented in Table III, although the text-only variant (w/ T5) achieves higher IS, the decline in fidelity metrics indicates that the text-only encoder fails to capture the intricate acoustic cues of the source audio. In contrast, Qwen2-Audio contributes richer cross-modal information by jointly encoding the instruction and the source audio, enabling more accurate and perceptually coherent edits than a text-only encoder could provide.
TABLE III: Ablation study on the instruction encoder and generation backbone.